# Investigating the heterogenous effects of a massive content moderation intervention via Difference-in-Differences

### v3

@@ 5 Discussion
Ethical considerations. This research furthers our understanding of the effects of content moderation by highlighting the complexities of user responses to moderation efforts. These new insights can contribute to the development of more robust and targeted interventions for reducing online toxicity while minimizing unintended side effects. Our study also discusses the ethical trade-off between the common good and minority harm. Specifically, the dilemma faced by moderators who must choose whether to enforce actions that, while benefitting the broader community, may inadvertently cause harm to a minority of users by increasing their resentment.
@@ 6 Conclusions
The Great Ban was a large-scale deplatforming intervention aimed at shutting down toxic communities on Reddit. To assess its effectiveness, we analyzed 53M comments made over 14 months by almost 34K users involved in the ban. Our findings show that 15.6% of the affected users left Reddit following the ban, while those who remained reduced their toxicity by an average of 4.1%. Despite this modest reduction in overall toxicity, 5% of users increased their toxicity by more than 70% compared to their levels before the ban. The impact of these resentful users was analyzed across the subreddits. We found that only 16% of users increased both their toxicity and activity and only 21% also received positive engagement from the community.
In general, in this study we provide new and more detailed insights into the effectiveness of The Great Ban, as well as its unintended consequences, Our findings can be beneficial to online platforms in the development of new moderation interventions and policies. For example, in future works we could investigate the relationship between users’ characteristics and the effects of a moderation intervention. This could lead to the development of targeted or personalized approaches that can be more effective in limiting the negative effects of moderation actions, such as the ones found in our study. Moreover, an encouraging direction for future research is the implementation of predictive models that forecast the effects of moderation interventions. Such models would be beneficial to content moderation, as they would allow moderators to plan and assess the effectiveness of certain moderation strategies.
@@ Acknowledgments
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity), and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).
Use of AI
During the preparation of this work the author(s) used ChatGPT in order to proofread the article prior to its submission. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.

### v4

@@ 5.6 Ethical considerations
This research furthers our understanding of the effects of content moderation by highlighting the complexities of user responses to moderation efforts. These new insights can contribute to the development of more robust and targeted interventions for reducing online toxicity while minimizing unintended side effects. Our study also discusses the ethical trade-off between the common good and minority harm. Specifically, the dilemma faced by moderators who must choose whether to enforce actions that, while benefitting the broader community, may inadvertently cause harm to a minority of users by increasing their resentment.
@@ 6 Conclusions
The Great Ban was a large-scale deplatforming intervention aimed at shutting down toxic communities on Reddit. To assess its effectiveness, we analyzed 53M comments made over 14 months by almost 34K users involved in the ban. Our findings show that 15.6% of the affected users left Reddit following the ban, while those who remained reduced their toxicity by an average of 4.1%. Despite this modest reduction in overall toxicity, 5% of users increased their toxicity by more than 70% compared to their levels before the ban. While the presence of resentful users was pervasive in the analyzed subreddits, their impact in terms of activity and obtained engagement was limited. We found that only 16% of users increased both their toxicity and activity and only 21% also received positive engagement from the community.
Our results provide new and more detailed insights into the effectiveness of The Great Ban, as well as its unintended consequences. Our findings can be beneficial to platform administrators and policymakers in the development of new moderation interventions and regulatory policies. For example, future works could investigate the relationship between user characteristics and the effects of a moderation intervention. This could lead to the development of targeted or personalized approaches that can be more effective in limiting the negative effects of moderation actions, such as the ones found in our study. Furthermore, an encouraging direction for future research is the implementation of predictive models that forecast the effects of moderation interventions. Such models would be beneficial to content moderation, as they would allow moderators to plan and estimate the effectiveness of certain moderation strategies ahead of their deployment.
@@ Acknowledgments
This work was supported by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006); by the ERC project DEDUCE under grant #101113826; by the PNRR-M4C2 (PE00000013) “FAIR-Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded under Next Generation EU; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT 4o to proofread the article prior to its submission. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.